A table top measures 2 m 25 cm by 1 m 50 cm. What is the perimeter of the table top?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Length of table top = 2 m 25 cm = 2.25 m
Breadth of table top = 1 m 50 cm = 1.50 m
Perimeter of table top = 2 (Length + Breadth)
= 2 (2.25 + 1.50)
= 2 (3.75) = 2 × 3.75 = 7.5 m
∴The perimeter of the table top is 7.5 m

Thanks